OpenStudy (anonymous):

G(03.06) Which of the following will give you the incorrect slope?

OpenStudy (anonymous):

(03.06) Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of . y - 6 = (x + 3) y = x - 6 y + 3 = (x - 6) x - 2y = 12